Fermi National Laboratory


Accelerator Update for August 8, 2005 - August 10, 2005

Monday, August 8
The day shift began with Operations monitoring store 4320 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, with Booster suffering from RF problems, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations established beam to the Meson MTest experiment at 8:50 AM.

At 9:13 AM, Operations reported that they were operating Booster at only four turns due to RF stability issues. The Run Coordinator scheduled Tuesday as a maintenance day for Booster RF repairs. Operations was to keep running Booster as best as they could.

A Switchyard (SY) expert began a study at 10:01 AM.

At 12:08 PM, Operations reported that BRF9 had sparked more than 350 times since 8 AM.

At 12:40 PM, BRF16 tripped off. An operator called an expert who had the station back up by 1:20 PM.

At 12:44 PM, BRF17 tripped off. It reset okay.

At 1:05 PM, BRF15 & 16 tripped off. They reset okay.

At 1:18 PM BRF3 & 15 tripped off. They reset okay

AT 1:24 PM, BRF3 tripped off a couple of times. It reset okay.

At 1:39 PM, BRF16 tripped off. It reset okay.

At 4:58 PM, BRF16 tripped off. It reset okay.

At 6:44 PM, Operations noted a drop in Pbar's antiproton production. Operations contacted an expert for advice and continued to tune.

Operations terminated store 4320 at 7 PM.

At 7:09 PM, BRF16 tripped off. It reset okay.

Operations began preparing the CDF and D0 collision halls for access at 7:17 PM.

The High Voltage Coordinator reported at 7:53 PM that feeder 47 had been energized.

D0 came out of access at 9:24 PM, but they went back in access at 9:52 PM to reset a tripped power supply.

CDF came out of access at 9:56 PM.

D0 reported at 10:15 PM that their Muon system power supply has failed and must be replaced. They couldn't run without this system. The Run Coordinator allowed them to begin the replacement.

At 10:59 PM, the MCenter experiment reported that they were having problems with their detector and requested that their beam be stopped while they investigated.

Operations reported at 11:58 PM that operators had completed a Pbar tune up and the antiproton production was back to normal.

Tuesday, August 9
The midnight shift began with Operations monitoring the Recycler stash, with Pbar stacking, with D0 in access to replace a power supply, with MiniBooNE and NuMI taking beam, and with MCenter off due to detector problems.

At 12:12 AM, Operations reported that a Linac RF station (KRF5) had tripped off for the first of four times in twelve minutes on a LCW flow fault indication. It reset each time.

At 12:55 AM, D0 reported that they could not complete their power supply replacement. A connector broke and they didn't have a replacement.

Operations put the TeV into a dry squeeze at 1:13 AM.

Operations put the TeV into shot setup at 1:46 AM.

At 2:09 AM, KRF5 tripped off. Operators investigated and had the station back on within ten minutes.

At 2:11 AM, BRF16 tripped off and would not reset. Operations called in an expert.

Operations established store 4322 at 4:42 AM. D0 could not get their luminosity counter online. CDF measured the luminosity at 113.2E30. Operations expected the initial average to be around 105.0E30. Pbar resumed stacking.

At 4:33 AM, during shot setup, Operations reported that a Main Injector (MI) Lambertson (I:LAM62) had tripped off. Operations decided to leave it off and have it checked out later in the morning.

Operations turned off the Preaccelerator at 6:31 AM, in preparation for Booster RF work and other machine maintenance.

Operations established beam to the Linac dump at 12:39 PM.

Operations established beam to the Booster dump at 5:40 PM.

Pbar resumed stacking at 5:55 PM.

Operations established beam to NuMI at 8:04 PM.

At 8:51 PM, Operations reported that they couldn't turn on I:LAM62. An operator called in experts.

Operations established beam to MiniBooNE around 9:16 PM.

At 10:30 PM, Operations reported that a Debuncher septum (D:ISEP) wasn't putting out the correct current. An operator called in an expert.

Operations prepared the Recycler at 11:32 PM to stash antiprotons.

Wednesday, August 10
The midnight shift began with Operations monitoring store 4322 and the stack, with Recycler in shot setup, and with MiniBooNE and NuMI taking beam.

The Recycler stashed ended and Pbar resumed stacking at 1:24 AM.

MCenter resumed taking beam at 1:32 AM.

At 1:54 AM, Operations reported that the Pbar beam permit tripped off and wouldn't reset. The alarm system wasn't displaying alarms, which made it hard to discover what was keeping Pbar off. Operations called an expert who investigated and had the alarm system working by 2:34 AM. Operators then found the Pbar device that had tripped the permit.

At 4:36 AM, BRF15 tripped off. It reset okay.

At 4:54 AM, BRF16 tripped off. It reset okay.

At 5:54 AM, BRF16 tripped off. It reset okay.

AT 6:14 AM, BRF15 tripped off. It reset okay.

The Plan for Wednesday
The plans for today are to first allow TeV experts conduct an end-of-store study until 9:30 AM and then prepare CDF and D0 for an access. Cryo system technicians will replace the flywheels on two wet engines at TeV sectors E3 & 4. Operations will put the TeV into shot setup sometime around 2 PM.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/10/2005   email Fermilab